OpenStudy (anonymous):

the perimeter of a rectangle equals the sum of the lengths of its four sides. If the width of a rectangle is 3 inches and its perimeter is 36 inches, find the length of the rectangle

OpenStudy (anonymous):

@Jhannybean @zzr0ck3r @Isaiah.Feynman

OpenStudy (jhannybean):

 The perimeter of a rectangle is fund by the equation \(\large P = 2l +2w\) where \(l\) = length and \(w\) = width. Were given \(w=3\) but not the length, so we labeled it an unknown variable. We're alsogiven the length of the perimeter, \(P= 36 \ in\) Now we plug this all into the equation. \[\large 36 = 2x +2(3)\] We're going to be solving for the length, therefore we need to isolate the x on one side of the equation. \[\large 36 - 6 = 2x\]\[\large x = \frac{30}{2}\] Can you solve for x?

OpenStudy (anonymous):

15 @Jhannybean

OpenStudy (jhannybean):

Yes, good job. :)

OpenStudy (jhannybean):

So the length = 15 = x. We can also check to make sure the length and width gives us 36. \[\large 36 = 2(3) + 2(15)\]\[\large 36 = 6 + 30\]\[\large 36 = 36 \ \checkmark \]